# Pipeline Parallelism with Controllable Memory

### v2

@@ E.2 Lower Bound
The runtime of a schedule can be bounded by any dependency path. In Figure 13, we present a dependency path which is a non-trivial lower bound for most schedules after squeezing and reordering. Assuming the actual peak memory is $k\bar{m}$ ($k\leq 2p$) and $T_{F}=T_{B}=T_{W}=1$, then the runtime of this dependency path is at least $16n+6p-3k-$ 1. Empirically, we find that we can always get close to this lower bound in our adaptive scheduler.
@@ Appendix F Building Blocks of V-Min and V-Half for all values of $d$
To avoid collisions between the repeating building blocks, we’ll need a slightly different offsets for different number of devices $d$. We list the details in Table 8 and 9 while continue using the notation defined in A. We didn’t list the offsets related to W passes because they always fill in the empty grids left by F and B. Some samples can also be found at Figure 14.
@@
$0\leq i<d,0<j\leq d$ $\delta F^{0}_{i}$ $\delta F^{1}_{j}$ $\delta B^{0}_{i}$ $\delta B^{1}_{j}$ $\delta(F_{d}^{0},F_{d}^{1})$ $\delta(F_{0}^{1},B_{0}^{1})$ $\delta(B_{d}^{0},B_{d}^{1})$
$0\equiv d\mod 3$ 1 1 1 1 1 3 1
$0<d\mod 3$ 1 1 1 1 1 1 1
@@
$0\leq i<d,0<j\leq d$ $\delta F^{0}_{i}$ $\delta F^{1}_{j}$ $\delta B^{0}_{i}$ $\delta B^{1}_{j}$ $\delta(F_{d}^{0},F_{d}^{1})$ $\delta(F_{0}^{1},B_{0}^{1})$ $\delta(B_{d}^{0},B_{d}^{1})$
$0\equiv d\mod 2$ 2 1 2 1 2 4 1
$1\equiv d\mod 2$ 2 1 2 1 2 1 1
Appendix G A Gallery of Pipeline Parallel Schedules and Their Building Blocks
We show the building blocks and full schedules of some well-known exsiting methods in Figure 18.
Appendix H Non-uniform Repeating Interval of Interleaved 1F1B
While most existing schedule in Figure 18 is repeated with uniform interval, Interleaved 1F1B [Narayanan et al., 2021] is slightly different in the repeating pattern. In Figure 15 we show two different repeating intervals. Official Interleaved 1F1B has a repeating interval as shown in Figure 15(a). If we modify the repeating interval to Figure 15(b), we’ll obtain another schedule with the same memory and bubble rate as official Interleaved 1F1B, shown in the gallery 18(i).
(a) The Non-uniform repeating interval of Interleaved 1F1B. The number in highlighted grids shows the minibatch with that index starts at this cell
@@ Appendix H Non-uniform Repeating Interval of Interleaved 1F1B
Appendix I Other Memory Efficient Building Blocks
Under the guidance of lifespan defined in 2.2, we can also find some of other building blocks besides the V-Shape building block family. We show the building blocks in Figure 16 and their full schedule in Figure 17. Both these schedules have lower bubble rate than 1F1B.
@@ Appendix I Other Memory Efficient Building Blocks
Using the building block defined in Figure 16(c), we can have an schedule Figure 17(c) with the same bubble rate as Interleaved 1F1B but lower memory.

### v3

@@ E.2 Lower Bound
The runtime of a schedule can be bounded by any dependency path. In Figure 13, we present a dependency path which is a non-trivial lower bound for most schedules after squeezing and reordering. Assuming the actual peak memory is $km$ ($k\leq 2d$) and $T_{F}=T_{B}=T_{W}=1$, then the runtime of this dependency path is at least $16n+6d-3k-$ 1. Empirically, we find that we can always get close to this lower bound in our adaptive scheduler.
@@ Appendix F Building blocks of V-Min and V-Half for all values of $d$
To avoid collisions when repeating building blocks, we need a slightly different offsets for different number of devices $d$. We list the details in Table 8 and 9 while continue using the notation defined in A. We do not list the offsets related to W passes because they always fill in the empty grids left by F and B. Some samples can also be found at Figure 14.
@@
$1\leq i<d,1<j\leq d$ $\delta F^{0}_{i}$ $\delta F^{1}_{j}$ $\delta B^{0}_{i}$ $\delta B^{1}_{j}$ $\delta(F_{d}^{0},F_{d}^{1})$ $\delta(F_{1}^{1},B_{1}^{1})$ $\delta(B_{d}^{0},B_{d}^{1})$
$d\equiv 0\mod 3$ 1 1 1 1 1 3 1
$d\not\equiv 0\mod 3$ 1 1 1 1 1 1 1
@@
$1\leq i<d,1<j\leq d$ $\delta F^{0}_{i}$ $\delta F^{1}_{j}$ $\delta B^{0}_{i}$ $\delta B^{1}_{j}$ $\delta(F_{d}^{0},F_{d}^{1})$ $\delta(F_{1}^{1},B_{1}^{1})$ $\delta(B_{d}^{0},B_{d}^{1})$
$d\equiv 0\mod 2$ 2 1 2 1 2 4 1
$d\equiv 1\mod 2$ 2 1 2 1 2 1 1
Appendix G Non-uniform Repeating Interval of Interleaved 1F1B
While most existing schedules (Figure 18) are repeated with uniform interval, interleaved 1F1B [Narayanan et al., 2021] is slightly different in the repeating pattern. The official interleaved 1F1B has a repeating pattern as shown in Figure 15(a). If we also employ uniform repeating interval as in Figure 15(b), we can obtain another schedule with the same memory footprint and bubble rate as the official interleaved 1F1B, shown in the gallery (Figure 18(i)).
(a) The non-uniform repeating interval of interleaved 1F1B. The number in highlighted grids shows the microbatch with that index starts at this cell.
@@ Appendix G Non-uniform Repeating Interval of Interleaved 1F1B
Appendix H Other Memory Efficient Building Blocks
Under the guidance of lifespan defined in Section 2.2, we can also find some other building blocks besides the V-Shape building block family. We show the building blocks in Figure 16 and their full schedules in Figure 17. All these schedules have lower bubble rate than 1F1B.
@@ Appendix H Other Memory Efficient Building Blocks
Using the building block defined in Figure 16(c), we can make a schedule with the same bubble rate as interleaved 1F1B but lower memory, shown in Figure 17(c).
@@
Appendix I A Gallery of Pipeline Parallel Schedules and Their Building Blocks
We show the building blocks and full schedules of some well-known existing methods in Figure 18.